I don't know how to solve a problem like this
9^2 + b^2 = 8^2
How do i solve for the b? Do i take the square root of c? or do i square it?

Answer by stanbon(75887)

9^2 + b^2 = 8^2
How do i solve for the b?
----------------
b^2 = 8^2 - 9^2
---
b^2 = 64 - 81
b^2 = -17
----
b^2+17 = 0
(b+isqrt(17))(b-isqrt(17)) = 0
b = -isqrt(17) or b = isqrt(17)
=================================
Cheers,
Stan H.